OpenStudy (anonymous):

What is the 25th term of the arithmetic sequence 7, 13, 19, … ?

OpenStudy (callisto):

Common difference = 13-7 = 6 T(n) = a + (n-1)d T(25) = 7 + (25-1) (6) = ...?